OpenStudy (anonymous):

What is the solution to the equation 2 ^(1 + 2x) = 3 ? A) x ? -0.292 B) x ? -0.185 C) x ? 0.185 D) x ? 0.292

OpenStudy (anonymous):

x=-9

OpenStudy (zepp):

The first thing you have to do here is to move the 21 to the left side, to do this, we can substract 21 to all sides \[21+2x-21 = 3-21\]\[2x = -14\]

OpenStudy (anonymous):

2x would equal -18

OpenStudy (zepp):

Then divide by 2 to all sides, \[2x = -18\]\[2x \div 2 = -18 \div 2\]\[x = -9\] Yeah right, -18, not -14, what was I thinking..